one number is five more than another. their sum is -17. fin the numbers
Answer by ikleyn(53937)

.
1.   One equation setup and solution

    x + (x-5) = -17


    2x + 5 = -17  ====>  2x = -17-5 = -22  ====>  x = -11.


    Answer.  The numbers  are  -11  and  -11+5 = -6.


2.   Two equations setup and solution

    x + y = -17     (1)
    x - y =   5     (2)
    -------------------------Add the equations to eliminate y. You will get


    2x = -17 + 5 = -12   ====>  x = -6


    Then from  (1)   y = -17 - x = -17 - (-6) = -17 + 6 = -11,


    and you get the same answer:  The numbers  are  -11  and  -6.